OpenStudy (pumpkin1102):

Josh and his friends bought vanilla wafers for $4 per packet and chocolate wafers for $1 per packet at a carnival. They spent a total of $45 to buy a total of 27 packets of wafers of the two varieties.

OpenStudy (pumpkin1102):

How many packets of chocolate wafers and vanilla wafers did they buy? Explain how you got the answer and why you selected a particular method to get the answer.

OpenStudy (erak):

If I set up the two equations would you be able to solve it on your own?

OpenStudy (pumpkin1102):

Maybe I would need to see first I am a visual person

OpenStudy (erak):

Okay so you're gonna have two variables, let's use x for vanilla, y for chocolate. This equation is the amount of wafers: x + y = 27 Where some amount of vanilla + chocolate is 27 wafers in total. The next equation is: 4x + y = 45 This is your dollar equation, because vanilla is $4 per packet and chocolate is $1, and they add up to $45.

OpenStudy (pumpkin1102):

😁

OpenStudy (pumpkin1102):

i don't get it

OpenStudy (erak):

Which part? Do you get that x + y = 27 is measured in amount of packets, and 4x + y = 45 is all money?

OpenStudy (pumpkin1102):

yes but, why?

OpenStudy (erak):

it's the information you are given and it's the only way you're going to relate money + amounts of wafers. You basically take those two and manipulate them for a variable and substitute them into the other.

OpenStudy (pumpkin1102):

so I use those 2 equations and combine them to make 1 equation?

OpenStudy (erak):

Well would you like me to show you how to do it and you try to follow along?

OpenStudy (pumpkin1102):

yes

OpenStudy (erak):

When given: x + y = 27 and 4x + y = 45 Solve equation 2 for either variable, doesn't really matter, I'm going with y. 4x + y = 45 y = 45 - 4x Next, plug this new y value into equation 1 and simplify x + y = 27 x + 45 - 4x = 27 3x = 18 x = 6 Now that you know x = 6, plug back into either equation to get y. x + y = 27 6 + y = 27 y = 21

OpenStudy (pumpkin1102):

O I understand it that wasn't as hard as I thought it would be

OpenStudy (erak):

Yeah as long as you get it and can do questions similar to this then you're good

OpenStudy (pumpkin1102):

Yeah I had been doing that I guess it was just the question itself that sounded confusing